OpenStudy (anonymous):

Find a 3x1 matrix X with entries not all zero suchthatAX =3X , where A= (1,2 ,-1) (1,0 ,1) (4,-4,5)

OpenStudy (anonymous):

\[ㅇㄹㄴㅁㄴㅇㄹ\nu \theta \Gamma \Psi \mathbb{Z}^-\delta\]

OpenStudy (anonymous):

This is called an eigenvector problem. If you search on wikipedia or something you'll find some solutions. I haven't done these for a while, but I think usually they don't give you the 3... AX = 3X, then AX - 3X = 0 (A-3I)X = 0 I is the identity matrix, which is 3*3 in dimensions. Now you can work out the matrix A-3I. 3I = [3, 0, 0 0,3,0 0,0,3 ] So, calculate A-3I. Now you just have the type of problem which is BX = 0. Solve that in the normal way to find the vector X.

OpenStudy (experimentx):

trivial solution would be 3x1 zero matrix

OpenStudy (anonymous):

all solutions are multiples of [-1/4, 1/4, 1]. You can get that if you follow the method above.

OpenStudy (experimentx):

given matrix, eigen value, and find eigen vector huh??